A patient (63kg) is treated with an experimental drug for hypertension. The drug is available as tablets with oral bioavailability of 88%. If the Vd is 2L/kg and the desired concentration is 2.8 mg/L, which of the following is the most appropriate oral loading dose for this drug?

207 mg

400.9 mg

110mg

.4 mg

613mg
Answer #1

ANSWER.

* Calculation of loading dose.

      Formula

        Loading dose = Vd(Volume of distribution) * (Target concentration)Tc / F(Bioavailablity).

Here

      Volume of distribution = 2 L/kg

      Volume of distribution for patient weight 63 kg = 2 * 63

                                                                                = 126 L.

        Target concentration = 2.8 mg/L.

              Bioavailability = 88%

                                     = 88/100

                                     = 0.88.

As per above formula

            Loading dose = 126 * 2.8 / 0.88

                                   = 352.8 / 0.88

                                   = 400.9 mg.